# GenCO: Generating Diverse Solutions to Design Problems with Combinatorial Nature

### v1

@@ 5 Conclusion
The underlying idea of our framework is to combine the flexibility of deep generative models with the guarantees of optimization solvers. By training the generator end-to-end with a surrogate linear combinatorial solver, GenCO generates diverse and combinatorially feasible solutions, with the generative loss being computed only on feasible solutions.
We have tested GenCO on various combinatorial optimization problems and generative settings, including GAN in Zelda game level generation and Warcraft map generation for path planning, demonstrating the flexibility of our framework for integrating into different generative paradigms. Our framework consistently produced diverse and high-quality solutions that satisfy the combinatorial constraints, which can be flexibly encoded using general-purpose combinatorial solvers.
Appendix A Detailed experimental setup for map generation
@@ A.1 Pseudocode
1:  Initialize generator parameters $\theta_{gen}$;
2:  Initialize adversary parameters $\theta_{adv}$;
3:  Input: distribution of the true dataset ($c_{true}\sim p_{\text{data}(c)}$), GAN’s objective $\mathcal{L}(\theta_{gen},\theta_{adv})$;
@@
6:        Sample a noise $\epsilon$;
7:        Sample true examples from dataset $c_{true}\sim p_{\text{data}}(c)$;
8:        Sample fake examples using $c_{fake}\sim G(\epsilon;\theta_{gen})$;
9:        Transform $c_{fake}$ into the coefficients of the optimization problem: $c=g(c_{fake})$;
10:        Solve: $x^{*}=\operatorname*{arg\,min}_{x\in\Omega}c^{T}z$;
11:        Backpropagate $\nabla_{\theta_{gen}}\left[\mathcal{L}(c_{fake};\theta_{gen},\theta_{adv})+\beta c^{T}x^{*}\right]$ to update $\theta_{gen}$;
12:        Backpropagate $\nabla_{\theta_{adv}}\left[-\mathcal{L}(c_{fake};\theta_{gen},\theta_{adv})+\mathcal{L}(c_{real};\theta_{gen},\theta_{adv})\right]$ to update $\theta_{adv}$;
@@
Algorithm 3 provides a detailed description of the GenCO framework in our penalty formulation and utilized in section 4.2. In this process, we sample both real and synthetic data, drawing from the true data distribution and the generator $G$ respectively (lines 7–8). Subsequently, the synthetic data undergoes a fixed mapping (e.g. ResNet in our experiments), called cost neural net (or cost NN), to obtain coefficients for the optimization problem, specifically the edge weights for the Shortest Path. Following this, we invoke a solver that provides us with a solution and its associated objective (lines 10–11). We then proceed to update the parameters of the generator $G$ using both the GAN’s objective and the solver’s objective. Finally, we refine the parameters of the adversary (discriminator) in accordance with the standard GAN’s objective.
@@ A.2 Settings
We employ ResNet as the mapping $g(\cdot)$ from Algorithm 3, which transforms an image of a map into a $12\times 12$ grid representation of a weighted directed graph: $g:\Re^{96\times 96\times 3}\rightarrow\Re^{12\times 12}$. The first five layers of ResNet18 are pre-trained (75 epochs, Adam optimizer with lr=$5e-4$) using the dataset from Pogančić et al. (2020), comprising 10,000 labeled pairs of image–grid (refer to the dataset description below). Following pretraining, we feed the output into the Shortest Path solver, using the top-left point as the source and the bottom-right point as the destination. The resulting objective value from the Shortest Path corresponds to $f$.
@@
The dataset used for training in the Shortest Path problem with $k=12$ comprises 10,000 randomly generated terrain maps from the Warcraft II tileset Pogančić et al. (2020) (adapted from Guyomarch (2017)). These maps are represented on a $12\times 12$ grid, with each vertex denoting a terrain type and its associated fixed cost. For example, a mountain terrain may have a cost of 9, while a forest is assigned a cost of 1. It’s important to note that in the execution of Algorithm 3, we don’t directly utilize the actual (ground truth) costs, but rather rely on ResNet to generate them.
@@ A.3 Architecture
We employ similar DCGAN architecture taken from Zhang et al. (2020) (see fig. 3 therein). Input to generator is 128 dimensional vector sampled from Gaussian noise centered around 0 and with a std of 1. Generator consists of five (256–128–64–32–16) blocks of transposed convolutional layers, each with $3\times 3$ kernel sizes and batch normalization layers in between. Discriminator follows by mirroring the same architecture in reverse fashion. The discriminator mirrors this architecture in reverse order. The entire structure is trained using the WGAN algorithm, as described inZhang et al. (2020).
@@ Appendix B GANs with combinatorial constraints
$\mathcal{L}_{\text{gen}}{}(G(\epsilon;\theta_{gen}))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(G(\epsilon;\theta_{gen})))\right]$ (4)
@@
$\min_{\theta_{gen}}\quad$ $\mathcal{L}_{\text{gen}}{}(S(\tilde{G}(\epsilon;\theta_{gen})))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(S(\tilde{G}{}(\epsilon;\theta_{gen}))))\right]$ (5)
where $\tilde{G}{}$ is unconstrained generator and $S$ is a surrogate combinatorial solver as described above. Here, we also have adversary’s learnable parameters $\theta_{adv}$. However, that part does not depend on combinatorial solver and can be trained as in usual GAN’s. The algorithm is presented in pseudocode 4.
  Initialize generator parameters $\theta_{gen}$;
  Initialize adversary parameters $\theta_{adv}$;
@@
        Sample problem $\epsilon{}$;
        Sample true examples from dataset $x_{true}\sim p_{\text{data}}(x)$;
        Sample linear coefficients $c\sim G(\epsilon{};\theta_{gen})$;
        Solve $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$;
        Backpropagate $\nabla_{\theta_{gen}}\mathcal{L}_{\text{gen}}{}(x^{*};\theta_{gen})$ to update generator (equation 5);
        Backpropagate $\nabla_{\theta_{adv}}\left[\log(f_{\theta_{adv}}(x_{true}))-\mathcal{L}_{\text{gen}}{}(x^{*};\theta_{gen})\right]$ to update adversary;
@@
Algorithm 4 GenCO GenCO in the constrained generator setting
@@ Appendix C GenCO – VQVAE
The formulation below spells out the VQVAE training procedure. Here, we simply train VQVAE on a dataset of known objective coefficients, which solves the problem at hand. A variant of this also puts the decision-focused loss on the generated objective coefficients, running optimizer $g_{\text{solver}}$ on the objective coefficients to get a solution and then computing the objective value of the solution.
$L_{\text{ELBO}}(c,\theta,E)=\mathbb{E}_{q_{\theta}(z|c)}[\log p_{\theta}(c|z)]-\beta\cdot D_{\text{KL}}(q_{\theta}(z|c)||p(z))+\gamma\cdot\|sg(\mathbf{e}_{k})-\mathbf{z}_{e,\theta}\|_{2}^{2}$ (6)
Here $z$ is an embedding vector, $c$ is the objective coefficients, $\log p_{\theta}(c|z)$ is a loss calculated via the mean squared error between the decoder output and the original input objective coefficients, $q_{\theta}(z|c)$ is the encoder, $p(z)$ is the prior, and $sg(\cdot)$ is the stop gradient operator, $E$ is a discrete codebook that is used to quantize the embedding.
$\mathcal{L}_{\text{optimization}}=\mathbb{E}_{c\sim p_{\theta}(c\mid z)}\left[f_{\text{obj}}(g_{\text{solver}}(c;y))\right]$ (7)
@@
Input:Training data distribution $D$ over problem info $y$ and known high-quality objective coefficients $c$, regularization weight $\beta$, linear surrogate solver $g_{\text{solver}}$, nonlinear objective $f_{\text{objective}}$. Output:Trained encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and codebook $E$ Initialize the parameters of the encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and the codebook $E=\{e_{1},e_{2},\ldots,e_{K}\}$ with $K$ embedding vectors;
for $t=1$ to $T$ do
       Sample $y,c$ from the distribution $D$;
       Compute the encoder output $z_{e}=f_{\text{enc}}(y,c)$;
       Find the nearest embedding vector $z_{q}=\arg\min_{e\in E}\|z_{e}-e\|_{2}^{2}$;
       Compute the quantization loss $L_{\text{quant}}=\|z_{e}-z_{q}\|_{2}^{2}$;
       Compute the reconstruction $\tilde{c}=f_{\text{dec}}(y,z_{q})$;
       Compute the reconstruction loss $L_{\text{recon}}=\|c-\tilde{c}\|_{2}^{2}$;
       GenCo Variant: Compute the optimization loss $L_{\text{opt}}=f_{\text{objective}}(g_{\text{solver}}(\tilde{c},y))$;
       Compute the total loss: $L_{\text{total}}=L_{\text{recon}}+\beta_{1}L_{\text{quant}}$ + $\beta_{2}L_{\text{opt}}$;
       Update the parameters of the encoder, decoder, and codebook to minimize $L_{\text{total}}$;
end for

### v2

@@ 5 Conclusion
The underlying idea of our framework is to combine the flexibility of deep generative models with the guarantees of optimization solvers. By training the generator end-to-end with a combinatorial solver, GenCO generates diverse and combinatorially feasible solutions, with the generative loss being computed only on feasible solutions.
Currently, many of GenCO’s requirements are tied to those of the space of differentiable solvers which is a rapidly evolving field. Some of the key bottlenecks researchers are working to address are scalability, differentiability, and the range of optimization problems that can be differentiated. Another frontier of differentiable solver research is in learning constraints from data. Currently it is assumed that practitioners have access to a mathematical representation of the constraints that can be optimized over. However, in some domains such as molecular design, the constraints such as binding affinity are not available analytically and it is nontrivial to guarantee feasibility using mathematical programs. Learning mathematical representations of these constraints from data is an active area of research.
Additionally, as with many generative models, it is important to keep real-world deployment in mind. Even though GenCO can guarantee feasibility with respect to explicitly encoded constraints, it cannot yet guarantee feasibility with respect to constraints that are desirable but not explicitly encoded. For instance, unless explicitly encoded or leveraging other alignment research, GenCO cannot guarantee fairness, privacy, or positive social impact. Similarly, in image generation, it may be difficult to encode some constraints such as ensuring that pixels representing hands have exactly 5 fingers using a diffusion model in image generation. Overall, GenCO is a step towards enforcing desirable constraints on the outputs of generative models, but it does not immediately solve all problems in ensuring the safety of generative models.
Overall, we propose GenCO, the first generative model that generates diverse solutions, optimizing differentiable nonlinear objectives with guaranteed feasibility for explicitly combinatorial constraints. We test GenCO on three different scenarios: inverse photonic design, game level design, and path planning, showing that GenCO outperforms existing methods along multiple axes. We have tested GenCO on various combinatorial optimization problems and generative settings, including GAN in Zelda game level generation, Warcraft map generation for path planning, and inverse photonic design. These settings require that GenCO handle a variety of generative architectures (GAN / VAE), optimization problems (linear programs, quadratic programs, mixed integer programs), and solvers (SCIP, Gurobi, shortest path solvers, blackbox domain specific solvers), demonstrating the flexibility of our framework for integrating different generative paradigms and optimization methods. Our framework consistently produces diverse and high-quality solutions that satisfy combinatorial constraints, which can be flexibly encoded using general-purpose or domain specific combinatorial solvers.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning, deep generative models, and combinatorial optimization. There are many potential societal consequences of the fields, none which we feel must be specifically highlighted here. Additionally, while our work focuses on guaranteeing the output of generative models satisfies specified constraints, it may be difficult to characterize all types of constraints which may be impossible to specify or enumerate.
Acknowledgment
The research at the University of Southern California was supported by the National Science Foundation (NSF) under grant number 2112533. We also thank the anonymous reviewers for helpful feedback.
Appendix A Detailed experimental setup for map generation (section 4.3)
@@ A.1 Pseudocode
1:  Initialize generator parameters $\theta_{gen}$
2:  Initialize adversary parameters $\theta_{adv}$
3:  Input: distribution of the true dataset (${\bm{c}}_{true}\sim p_{\text{data}({\bm{c}})}$), GAN’s objective $\mathcal{L}(\theta_{gen},\theta_{adv})$
@@
6:        Sample a noise $\epsilon$
7:        Sample true examples from dataset $c_{true}\sim p_{\text{data}}(c)$
8:        Sample fake examples using ${\bm{c}}_{fake}\sim G(\epsilon;\theta_{gen})$
9:        Transform ${\bm{c}}_{fake}$ into the coefficients of the optimization problem: ${\bm{c}}=f({\bm{c}}_{fake})$
10:        Solve: $x^{*}={\bm{g}}({\bm{c}})=\operatorname*{arg\,min}_{x\in\Omega}{\bm{c}}^{T}x$
11:        Backpropagate $\nabla_{\theta_{gen}}\left[\mathcal{L}({\bm{c}}_{fake};\theta_{gen},\theta_{adv})+\beta c^{T}x^{*}\right]$ to update $\theta_{gen}$
12:        Backpropagate $\nabla_{\theta_{adv}}\left[-\mathcal{L}({\bm{c}}_{fake};\theta_{gen},\theta_{adv})+\mathcal{L}({\bm{c}}_{real};\theta_{gen},\theta_{adv})\right]$ to update $\theta_{adv}$
@@
Algorithm 2 provides a detailed description of the GenCO framework in our penalty formulation and utilized in section 4.3. In this process, we sample both real and synthetic data, drawing from the true data distribution and the generator $G$ respectively (lines 7–8). Subsequently, the synthetic data undergoes a fixed mapping (e.g. ResNet in our experiments), called cost neural net (or cost NN), to obtain coefficients for the optimization problem (the edge weights for the Shortest Path). Following this, we invoke a solver that provides us with a solution and its associated objective (lines 10–11). We then proceed to update the parameters of the generator $G$ using both the GAN’s objective (group loss $\mathcal{L}$) and the solver’s objective (individual loss $\mathcal{D}$). Finally, we refine the parameters of the adversary (discriminator) in accordance with the standard GAN’s objective.
@@ A.2 Settings
We employ ResNet as the mapping $f(\cdot)$ from the above Algorithm 2, which transforms an image of a map into a $12\times 12$ grid representation of a weighted directed graph: $f:\Re^{96\times 96\times 3}\rightarrow\Re^{12\times 12}$. The first five layers of ResNet18 are pre-trained (75 epochs, Adam optimizer with lr=$5e-4$) using the dataset from (Pogančić et al., 2020), comprising 10,000 labeled pairs of image–grid (refer to the dataset description below). Following pretraining, we feed the output into the Shortest Path solver, using the top-left point as the source and the bottom-right point as the destination. The resulting objective value from the Shortest Path corresponds to ${\bm{g}}$.
@@
The dataset used for training in the Shortest Path problem with $k=12$ comprises 10,000 randomly generated terrain maps from the Warcraft II tileset (Pogančić et al., 2020) (adapted from (Guyomarch, 2017)). These maps are represented on a $12\times 12$ grid, with each vertex denoting a terrain type and its associated fixed cost. For example, a mountain terrain may have a cost of 9, while a forest is assigned a cost of 1. It’s important to note that in the execution of Algorithm 2, we don’t directly utilize the actual (ground truth) costs, but rather rely on ResNet to generate them.
@@ A.3 Architecture
We employ similar DCGAN architecture taken from (Zhang et al., 2020) (see fig. 3 therein). Input to generator is 128 dimensional vector sampled from Gaussian noise centered around 0 and with a std of 1. Generator consists of five (256–128–64–32–16) blocks of transposed convolutional layers, each with $3\times 3$ kernel sizes and batch normalization layers in between. Discriminator follows by mirroring the same architecture in reverse fashion. The discriminator mirrors this architecture in reverse order. The entire structure is trained using the WGAN algorithm, as described in(Zhang et al., 2020).
@@ Appendix B GANs with combinatorial constraints
$\mathcal{L}{}(G(\epsilon;\theta_{gen}))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(G(\epsilon;\theta_{gen})))\right]$ (5)
@@
$\min_{\theta_{gen}}$ $\mathcal{L}{}(S(\tilde{G}(\epsilon;\theta_{gen})))=\mathbb{E}_{\epsilon{}}\left[\log(1-f_{\theta_{adv}}(S(\tilde{G}{}(\epsilon;\theta_{gen}))))\right]$ (6)
where $\tilde{G}{}$ is unconstrained generator and $S$ is a surrogate combinatorial solver as described above. Here, we also have adversary’s learnable parameters $\theta_{adv}$. However, that part does not depend on combinatorial solver and can be trained as in usual GAN’s. The algorithm is presented in pseudocode 3.
  Initialize generator parameters $\theta_{gen}$
  Initialize adversary parameters $\theta_{adv}$
@@
        Sample problem $\epsilon{}$
        Sample true examples from dataset $x_{true}\sim p_{\text{data}}(x)$
        Sample linear coefficients $c\sim G(\epsilon{};\theta_{gen})$
        Solve $x^{*}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$
        Backpropagate $\nabla_{\theta_{gen}}\mathcal{L}{}(x^{*};\theta_{gen})$ to update generator (equation 6)
        Backpropagate $\nabla_{\theta_{adv}}\left[\log(f_{\theta_{adv}}(x_{true}))-\mathcal{L}{}(x^{*};\theta_{gen})\right]$ to update adversary
@@
Algorithm 3 GenCO in the constrained generator setting
@@ Appendix C GenCO – VQVAE
The formulation below spells out the VQVAE training procedure. Here, we simply train VQVAE on a dataset of known objective coefficients, which solves the problem at hand. A variant of this also puts the decision-focused loss on the generated objective coefficients, running optimizer $g$ on the objective coefficients to get a solution and then computing the objective value of the solution.
$\mathcal{L}(\mathcal{X})=\text{ELBO}({\bm{x}},\theta,E)=\mathbb{E}_{q_{\theta}(z|{\bm{x}})}[\log p_{\theta}({\bm{x}}|z)]-\beta\cdot D_{\text{KL}}(q_{\theta}(z|{\bm{x}})||p(z))+\gamma\cdot\|sg(\mathbf{e}_{k})-\mathbf{z}_{e,\theta}\|_{2}^{2}$ (7)
Here $z$ is an embedding vector, $c$ is the objective coefficients, $\log p_{\theta}({\bm{x}}|z)$ is a loss calculated via the mean squared error between the decoder output and the original input objective coefficients, $q_{\theta}(z|{\bm{x}})$ is the encoder, $p(z)$ is the prior, and $sg(\cdot)$ is the stop gradient operator, $E$ is a discrete codebook that is used to quantize the embedding.
$\mathcal{D}=\mathbb{E}_{c\sim p_{\theta}(c\mid z)}\left[f_{\text{obj}}(g(c;y))\right]$ (8)
@@
1:  Given: Training data distribution $D$ over problem info $y$ and known high-quality solutions $x$, regularization weight $\beta$, linear surrogate solver $g_{\text{solver}}$, nonlinear objective $f_{\text{objective}}$.
2:  Output: Trained encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and codebook $E$
3:  Initialize the parameters of the encoder $f_{\text{enc}}$, decoder $f_{\text{dec}}$, and the codebook $E=\{e_{1},e_{2},\ldots,e_{K}\}$ with $K$ embedding vectors
4:  for $t=1$ to $T$ do
5:     Sample $y,x$ from the distribution $D$
6:     Compute the encoder output $z_{e}=f_{\text{enc}}(y,x)$
7:     Find the nearest embedding vector $z_{q}=\arg\min_{e\in E}\|z_{e}-e\|_{2}^{2}$
8:     Compute the quantization loss $L_{\text{quant}}=\|z_{e}-z_{q}\|_{2}^{2}$
9:     Decode the embedding $\tilde{c}=f_{\text{dec}}(y,z_{q})$
10:     Solve $\tilde{x}=\operatorname*{arg\,max}_{x\in\Omega}c^{T}x$
11:     Compute the reconstruction loss $L_{\text{recon}}=\|x-\tilde{x}\|_{2}^{2}$
12:     Compute the optimization loss $L_{\text{opt}}=f_{\text{objective}}(\tilde{x})$
13:     Compute the total loss: $L_{\text{total}}=L_{\text{recon}}+\beta_{1}L_{\text{quant}}+\beta_{2}L_{\text{opt}}$
14:     Update the parameters of the encoder, decoder, and codebook to minimize $L_{\text{total}}$
15:  end for